If a =/ 0 and the sum of x and 1/a is 0, then. (A) x = -1/a (B) x = -a (C) x = 1-a (D) x = a
Answer by stanbon(75887)

If a =/ 0 and the sum of x and 1/a is 0, then.
---
x + (1/a) = 0
x = -1/a
------------------
Cheers,
Stan H.
==============
(A) x = -1/a
(B) x = -a
(C) x = 1-a
(D) x = a
================================